Actions
2015-03-10 physics coordinator report » History » Revision 19
« Previous |
Revision 19/20
(diff)
| Next »
Rogers, Chris, 10 March 2015 14:57
physics coordinator report¶
- Step I papers:-
- (Polarisation analysis - held up by dodgy EMR code)
- Cooling demo paper:
- Secondary absorber did not present sufficient material to the beamline to
shield the tracker, review of secondary absorber position which has caused
delay to the paper. - Tried for a revision this week for review, it didn't work out, hope to have a
revision by next week
- Secondary absorber did not present sufficient material to the beamline to
- Step IV papers:
- Status at http://micewww.pp.rl.ac.uk/analysis/analysis_dashboard.html
- Trying to bring material physics people "out of the woodwork", conflict that Warwick + Glasgow both want to do MCS, no one wants dE/dx
- Rogers is re-establishing selection/weighting algorithms (dead since 2007)
- aim to drive an analysis of "physics block challenge" before MPB
- feed into review of data rate issues, etc.
- Field mapping
- Blackmore/Cobb have a good analysis of the field axis in the FC, which does not
produce the "tails" ruining the fitting, hopeful that SS will follow in time
for alignment beginning of April
- Blackmore/Cobb have a good analysis of the field axis in the FC, which does not
- March/April running
- Rogers will be physics shifter on March 21st
- Trying to establish physics shifters on March 28th and following
- June running
- Planning note in preparation, due Friday
- Discuss/finalise note at a meeting Thursday 19th
- Tracker will not be able to get plan in time for the meeting, rogers to phone Melissa
- EMR request one week data and additionally some cosmics
- C+M request one week
- TOF request calibration (1 or 2 shifts)
- KL request cosmic data taking
- Ckov request 1 shift
- Beamline (hardware) request an activation run
- Beamline (optics) not sure
- Cooling channel first data may at some point come through, not sure
Updated by Rogers, Chris over 8 years ago · 19 revisions